OpenStudy (anonymous):

What type of triangle is formed by the points A(4, 2 ), B(6,−1), and C(−1, 3 )? A right B equilateral C isosceles D scalene

OpenStudy (anonymous):

D is what i would say

OpenStudy (p0sitr0n):

AB = 13^1/2 BC = 65^1/2 AC = 26^1/2 not rectangle, not equilateral and not isosceles must be d?

OpenStudy (anonymous):

ok thanks everyone! :)

OpenStudy (anonymous):

D^2=(X2-X1)^2+(Y2-Y1)^2